cirle O centered at the orgin passes through points A(-4,3) and B(3,4). Use algebra to show that the perpendicular bisector of line AB passes through the center of the circle.
Answer by jedidiah(6)

y=mx+b
AB: m= 1/7
y=(1/7)x + b
4= 3/7 + b
b=25/7
y= (1/7)x + (25/7)
slope of line perpendicular to AB = -7
midpoint of AB = (-1/2, 7/2)
equation of perpendicular bisector:
y= -7x + b
7/2 = 7/2 + b
b=0
y=-7x
0=0